OpenStudy (anonymous):

how do you find minimum and maximum values for example, of y=2x^2-8x+5 ?

OpenStudy (anonymous):

in the quadratic equation \(\large y=f(x)=ax^2+bx+c \), the parabola will have a maximum if a<0, and have a minimum if a>0. the maximum/minimum is obtained by plugging in the value \(\large p=-\frac{b}{2a} \) into the function. the vertex is (p, f(p))

OpenStudy (anonymous):

would p be y or x

OpenStudy (anonymous):

p is the x-value

OpenStudy (anonymous):

ok so for this one is it x=2?

OpenStudy (anonymous):

yes.... that's the x-coordinate of the vertex... is the vertex a max or min???

OpenStudy (anonymous):

maximum

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what is the "a" in your equation? is a<0 ???

OpenStudy (anonymous):

a is 2 and it is not less than zero so its a minimum right ?

OpenStudy (anonymous):

yes... :)

OpenStudy (anonymous):

sorry got it confused

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

wait... one more thing...

OpenStudy (anonymous):

the question asks "how do you find the maximum/mininum". i'm assuming they want the maximum/minimum so you'll need to find f(2).

OpenStudy (anonymous):

could u explain this to me ?